Systems Staff Meeting Notes 10/16/00

Systems Staff Meeting Notes
October 16, 2000

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Debra Harris, Martha Hruska, Bobbie Meng, Suzy Shaw, Hilton Soto, Geoffrey Stewart, Ying Tang
Announcements
  1. Security System - camera will be installed out in Systems loading dock area. Motion detecting cameras, card swipes & surveillance cameras are all forthcoming.
  2. Gators won this weekend.
  3. Geoffrey cautions that there is a nasty flu bug going around.
Review Minutes of 10/9/00
  1. Announcements #6 - Gulf Coast University, not Gulf Coast Community College.
  2. RIR #7 - Server for Ref Chat - no date or timeframe given, but we need to go ahead and order a server. Two servers will not be necessary since they will be accessing via laptop. Will shall work on ordering a new one.
  3. Bobbie, Suzy & Will completed their MSSE training classes.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Digital said the new fiber switch will be here today.
  2. The 17" monitors have not been received yet.
  3. A distribution list as well as a list of old machines to be returned has been generated. There are still old Macs in Maps, Documents & HS&S.
  4. Six contact barcode scanners were received on Thursday. One was distributed to West Circulation Desk, one to Documents, two to MSL, one to Music and one to AFA. Will have each department test them out for two weeks and provide feedback on their performance. Bobbie tested them and they scanned consistently. If they work well, we will need to order quite a few more of them.
  5. One laptop we have on hand is for Reference Chat. One laptop is for Preservation and one for Resource Services. Will notes that we have two laptops in the vault that are to go with the wireless installation - Resource Services could use one of those. Martha agrees - we should issue both of the wireless laptops to Resource Services & they can use them without the wireless. When the wireless system comes to fruition we can change the card for the wireless setup. This would leave one laptop still available.
  6. We are out of 3 foot patch cables. Need to order more. Also need to order 2 foot ones as well. Variety of color would be good.
  7. Since we are going to be buying a server for Net Agent, what about a replacement for NT2? Should we order a new one - it is near full right now. Yes. The DLC server that is in the vault - Power Edge 4400 - we could build it into an Exchange server and transfer the data over. Martha will check with Erich on the degree of urgency for their server. We could give the old NT2 to Erich as the DLC server. Will needs to check the specs of the 4400 to see if it meets the requirements that Bill desires for the Exchange server.
Liaisons
  1. Geoffrey, Will and Bobbie have been collaborating on the training session. Geoffrey is developing topical handouts for the various application installations: McAfee, IE, Office, Hummingbird. Will also include information on how to configure IP's. Geoffrey will forward the information to Debra to update the liaison section of the Systems Web page.
  2. Geoffrey will conduct a dry run of the training session to Systems staff on Wednesday in the East conference room. The real training will probably run about 2 hours, with a 15 minute break followed by a Q&A session.
  3. The specific day of the training session is still up in the air. The list of attendees should be generated by Middle Managers. May give an outline of the session to the few super-liaisons for comment.
Networking
  1. DLC - Geoffrey is working to reconfigure their NT machines to Win'98. Bobbie notes that their electrical needs on the 2nd floor have not been fully addressed. The way the pods have been configured will require very long (and out-of-spec) patch cords. There are more units then power connections. The furniture & pods are in place - they may be moving down sometime this week. Facilities should be moving them - we'll call to see if a date has been set.
  2. Public printing - no developments.
  3. Proxy server - Bill came in and made some adjustments on Friday. Will has been tinkering with Linux.
  4. Cataloger's Desktop - Will needs to start trying to install it.
  5. Rewiring in room 200 is complete. Room 300 is on hold for the phones. Two additional pods have been added - they are not wired. Will add them to the to the plan for Digital.
  6. South Tower - is not complete. Digital will return later this week, they have a new crew and are training people. They will set up a date & time to pull the wires in room 210 & 310. The Grand Reading Room is scheduled for 3 pulls - expertise will be required due to the cosmetics of the room.
  7. There are some more pulls needed in MSL, Education and Journalism. AFA pulls are complete.
  8. Fiber fed switch was installed in room 210 West this morning. Need to move ILL, Accounting and the Directors over to the new wire.
Software Projects Status
  1. Film Log - complete, waiting on testing feedback.
  2. Hilton got together with Bobbie last week and is working on some changes to the IR database.
  3. Ying met with Stephanie and Erich this morning. She demoed what she has & they have requested redesigns of the basic structure.
  4. Cynthia has modified the read-only IR database for Facilities - she'll add a wildcard capability for the search.
  5. MARCon/FULoader motif - Winston to modify the program (at least at the record preparation level) that automates the transfer of bibliographic records from LUIS to NOTIS and back. Martha will keep Winston informed.
  6. Library Web Group - co-chaired by Rich & Suzy - Martha used Outlook to send out tasks to Suzy who, in turn, sent them out to various committee members. It enables the supervisor to track things. The committee itself is trying to devise guideline documents for Web standards for the Libraries, trying to reach a greater consistency.
DP Coordinator
  1. The patron that Leilani was in contact with about purchasing one of our tapes has apparently declined the offer, since he has not returned her inquiries.
  2. Dolores was not too clear about the purpose of the inventory of tapes in the Tape Library that Debra sent to her. Martha will talk to Dolores and John Ingram.
Calendar
  1. Winston out Oct. 16, 19 & 20.
  2. Martha out Tues. Oct. 17 - Orlando - Tech Services.
  3. Ying out Oct. 18-20.
  4. Will and Hilton both out Nov. 20-22 & 27.


Back to the Systems Homepage

Last updated October 23, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu